# Both Ends Count! Just How Good are LLM Agents at Text-to-“Big SQL”?

### v2

@@ 2.1. Limitations of Current Metrics
The major limitation of these metrics is their reliance on all-or-nothing correctness [^2] The corresponding formulas are provided in the Appendix. And extended version of this work is available on arXiv: https://arxiv.org/abs/2602.21480. In practice, however, a generated query may deviate from the expected result without being entirely invalid, provided that end users may still be able to determine its correctness through simple validation, for example, by detecting an additional column in the projected output. We summarize the possible outcomes of an SQL execution as follows:
@@ 4.1. Accuracy is Not Enough in Text-to-Big SQL
Later-generation models do not clearly outperform their predecessors in zero-shot agentic text-to-SQL. For example, Opus 4.6 achieves perfect accuracy but takes 92.37% longer execution time than GPT-4o. Similarly, Gemini 3 Pro exhibits poor latency, further exacerbated by high variance from API instabilities. Notably, GPT-4o was released in 2024, nearly two years before the other two models.
@@ Acknowledgements.
This work has been partly funded by the EU Horizon programme under the grant agreements: 101092644 (NearData), 101092646 (CloudSkin), 101093110 (EXTRACT) and 101086248 (CLOUDSTARS). Germán T. Eizaguirre is recipient of a pre-doctoral FPU grant from the Spanish Ministry of Universities (ref. FPU21/00630). Marc Sánchez-Artigas is a Serra Húnter Fellow.
@@ Appendix A Agent Architecture
Figure 5. The architecture of our agent is centered on an executor built with LangGraph. For our evaluation, both the controller and the checker operate on the same LLM. The Spark Session can be connected to either a local deployment or a distributed cluster.
@@ D.2. Detailed Results
Query ID Model EX VES VES* VCES SVQ
@@ Appendix F Detailed Results
Table 5. Per-query TPC-H (SF1) metrics by model (mean $\pm$ std across runs). VES, VES*, and VCES follow the paper definitions; SVQ reports the expected cost per valid query.
Query ID Model EX VES VES* VCES SVQ

### v3

@@ 2.1. Limitations of Current Metrics
The major limitation of these metrics is their reliance on all-or-nothing correctness [^2] The corresponding formulas are provided in the Appendix. An extended version of this work is available on arXiv: https://arxiv.org/abs/2602.21480. In practice, however, a generated query may deviate from the expected result without being entirely invalid, provided that end users may still be able to determine its correctness through simple validation, for example, by detecting an additional column in the projected output. We summarize the possible outcomes of an SQL execution as follows:
@@ 4.1. Accuracy is Not Enough in Text-to-Big SQL
Later-generation models do not clearly outperform their predecessors in zero-shot agentic text-to-SQL. For example, Opus 4.6 achieves perfect accuracy but takes 92.37% longer execution time than GPT-4o. Similarly, Gemini 3 Pro and GLM-5 exhibit poor latency, further exacerbated by high variance from API instabilities. Notably, GPT-4o was released in 2024, nearly two years before the other two models.
@@ Acknowledgements.
This work has been partly funded by the EU Horizon programme, grant no. 101092646 (CloudSkin), and by the Spanish MICIU/AEI, grant no. PID2023-148202OB-C21. Germán T. Eizaguirre is recipient of a pre-doctoral FPU grant from the Spanish Ministry of Universities (ref. FPU21/00630).
@@ Appendix A Agent Architecture
The architecture of our agent is centered on an executor built with LangGraph, as illustrated in Figure 5. For our evaluation, both the controller and the checker operate on the same LLM, and the same controller is used for all iterations of the ReAct framework. LLMs could, however, be tuned for performance or cost-efficiency at the iteration level. While the tool call sequence in Figure 5 corresponds to the most common execution flow, this sequence may vary depending on the controller LLM (e.g., through tool repetition or omission). The Spark Session can be connected to either a local deployment or a distributed cluster. Although this work focuses on Spark, the text-to-Big SQL workflow is compatible with any data analytics framework and backend deployment.
Figure 5. Architecture and typical execution flow of the evaluated LLM agent.
@@ Appendix C Text-to-SQL formulas
Below we list standard formulas for text-to-SQL accuracy evaluation metrics, which serve as the foundation for the text-to-Big SQL metrics introduced in Section 2.
@@ C.1. Exact Matching (EM)
The $Q_{i}=\hat{Q}_{i}$ operation compares two queries to determine their equivalence. Query equivalence depends on context. Therefore, multiple methods have been developed to address it. A common metric is the Spider exact matching accuracy (yu-etal-2018-spider), which verifies whether generated SQL queries match gold-standard references in both structure and specific components. To achieve this, this metric decomposes a SQL query into its constituent clauses, such as SELECT, WHERE, HAVING, GROUP BY, and ORDER BY, and performs set-based comparisons. This approach effectively ignores ordering differences, such as varying sequences in conditions, to focus on semantic equivalence.
@@ C.2. Execution Accuracy (EA)
Usually, $V_{i}=\hat{V}_{i}$ uses a set-based equality: rows and columns must match exactly, ignoring order unless sorting is specified
@@ D.2. Detailed Results
Query ID Model Accuracy (Eq. 2) VES VES* VCES SVQ
@@ Appendix F Detailed Results
Table 5. Per-query average metrics in TPC-H (SF1) by model (mean $\pm$ std across runs). VES, VES*, and VCES follow the paper definitions; SVQ reports the expected cost per valid query.
Query ID Model Accuracy (Eq. 2) VES VES* VCES SVQ
@@
Appendix G Error Analysis
In the BIRD evaluation we executed 10 queries [^5] Query 886 was excluded from the main paper results since the Equation 7 results are low for most models. The query asks "Which year has the most number of races? The most number of races refers to max(round)"; most models returned the superfluous column max(round) instead of only the year. on 11 models [^6] DeepSeek Chat is omitted in the main paper results due to its comparatively poor performance. with 50 few-shot iterations of the translation for each, resulting in 5,500 total query executions. Out of these, 930 translations produced a result different from the expected output, yielding an Equation 7 result of 0. Table LABEL:tab:failed-iterations-per-model presents the per-model aggregate distribution of the errors.
Table 7. Total number of failed translations (Eq. 2==0==0) per model. Each model was evaluated on all 10 BIRD queries with 50 translations per query (500 iterations per model).
Model Failed Translations
DeepSeek Chat 214
GPT-5.2 190
GPT-5 104
GPT-4o 78
Gemini 2.5 Flash 70
Kimi K2.5 57
Gemini 3 Flash 51
GLM-5 51
Claude Opus 4.5 50
Claude Opus 4.6 50
Gemini 3 Pro 15
To provide a better analysis, we classify the obtained errors based on the taxonomy from (shen2025studyincontextlearningbasedtexttosqlerrors). We use their proposed categories because they also focus on LLMs for text-to-SQL translation and provide an extensive categorization of possible outcomes. Each query with a result different from the expected output (‘incorrect’ from now on) could be affected by one or more errors. We manually gathered 1,730 errors across the 930 incorrect queries. We depict the error taxonomy in Figure 7 and complement the diagram with a pie chart of the same data in Figure 8.
We extracted several interesting insights from the analysis. Within the Output Format (F2) errors, 627 (93.16%) were due to additional columns, meaning the result contained the expected data accompanied by unnecessary columns. Of those, 77.51% were due to just one additional column. Such results would be considered correct under our text-to-Big SQL metrics (Eq. 2). For instance, query 886 is considered incorrect by classical text-to-SQL metrics primarily because of an additional column (see Listing 2). We consider our metric to better represent valid queries, as a user could easily differentiate the required column in a practical case. We also identified an LLM difficulty in differentiating MAX and COUNT cases. Among Unaligned Aggregation Structure (E5) errors, 15.80% resulted from adding unnecessary MAX aggregations, while 5.41% were due to unnecessary COUNTs; notably, 44.16% involved mixing MAX and COUNT aggregators. Finally, in four queries, the agents ran an inspection operation (DESCRIBE table or SHOW TABLES) as the final query, a curious behavior despite the low number of cases.
Listing 1: Correct query for Q886
SELECT year
FROM races
GROUP BY year
ORDER BY COUNT(round) DESC
LIMIT 1
Listing 2: Incorrect query for Q886 (claude-opus-4-5)
SELECT year, MAX(round) as max_races
FROM races
GROUP BY year
ORDER BY max_races DESC
LIMIT 1
Figure 6. Correct and incorrect SQL examples for query 886.
We only tested the agent extensively on a subset of the BIRD benchmark; therefore, these results should not be considered absolute. However, we believe these insights provide valuable information for future research on text-to-Big SQL agents.
Error Taxonomy (1730 errors)Schema(5)Logic(79)Convention(20)Semantic(952)Not an Error(674) B1 Table-Column Mismatch (3) B2 Non-Existent Schema (2) C1 Implicit Type Conversion (79) D1 Violating Value Specification (20) E1 Incorrect Table Selection (149) E2 Projection Error (198) E4 Improper Condition (138) E5 Unaligned Aggregation Structure (462) E7 ORDER-BY Error (3) E9 Comparing Wrong Columns (2) F1 Gold Error (1) F2 Output Format (673)
Figure 7. Taxonomy (shen2025studyincontextlearningbasedtexttosqlerrors) of SQL translation errors from our LLM agent. Numbers in parentheses denote error counts. Categories: A=Syntax, B=Schema, C=Logic, D=Convention, E=Semantic, F=Not an Error. Total: 1730 errors across 5 categories.
Figure 8. Distribution of identified text-to-SQL translation errors across all 930 incorrect BIRD query translations, categorized by the taxonomy proposed in (shen2025studyincontextlearningbasedtexttosqlerrors)